OpenStudy (jozelynw):

The rhombus below has vertex coordinates shown. Find the coordinates of vertex S

OpenStudy (mathmale):

Please share an image of your rhombus. Write in whichever coordinates of the vertex are given. You may be able to figure out the unknown coordinates just by looking at your diagram. If not, refresh your memory regarding what a "rhombus" is.

OpenStudy (mathmale):

@JozelynW : OpenStudy says you're "just looking around." That's OK, but you need to consider what that tells me about your interest in solving this problem.

OpenStudy (fixxer):

-1,3

OpenStudy (fixxer):

means its 1 to the left of origin and 3 above the origin

OpenStudy (jozelynw):

There's no equation to find it..because I feel it couldn't have been that simple?

OpenStudy (fixxer):

Not that I know of, it is simple to see it graphical, that should be a totally fine answer :) Dont make stuff more advanced then necesarry is the key to being good in math i think :)

OpenStudy (jozelynw):

ok thanks

OpenStudy (fixxer):

You can draw the grids if it is for homework to show how u can see it :)

OpenStudy (mathmale):

Hint: Side RS is parallel to and has the same length as side TU.

OpenStudy (jozelynw):

so what would you do to find the S then?

OpenStudy (jozelynw):

I just want to know what does that have to do with the answer... How am I suppose to use the information you just gave me

OpenStudy (fixxer):

use the slope of the line, make them equal in length, either way you solve it visually, it is not a "equation" either way, so why make a complicated vizualization? Just my thoughts tho, as else method aint specified. if it said find S from the slope of RS or ie i would do it

OpenStudy (fixxer):

R=(-4,-1) the line goes up 4 and right 3 that can you see from the other paralell line U=0,-4 T=3,0 hence S=P+(3,4)=(-1,3)

OpenStudy (mathmale):

Actually, you may be able to read the answers (coordinates of point S) directly from the graph. What is the x-coord. of point S? the y-coord.? Take a close look.